### v2

# A note on Shintani’s invariant

Abstract.
Shintani’s celebrated invariants are conjectured to generate abelian extensions of real quadratic number fields, offering a potential solution to Hilbert’s 12th problem in that setting. In this note, we derive new expressions for Shintani’s invariants by generalizing an observation of Yamamoto, who showed that these invariants - originally formulated using the double sine function - can be expressed in terms of the q-Pochhammer symbol.
1. Introduction
In his seminal paper [7], Shintani formulated a conjecture stating that certain invariants
(1.1) $X(\mathfrak{f})=\prod_{k=0}^{g(\mathfrak{f})}\mathcal{S}(\varepsilon,z_{k})\mathcal{S}(\varepsilon^{\prime},z_{k}^{\prime}),$
defined via products of special values of the double sine function $\mathcal{S}(\omega,z)$, generate abelian extensions of real quadratic number fields. If true, this conjecture would provide a solution to Hilbert’s 12th problem in the real quadratic setting. However, the conjecture remains completely open, due to the absence of a geometric framework analogous to that of elliptic curves with complex multiplication, which plays a central role in the classical theory for imaginary quadratic number fields.
Despite extensive and convincing numerical evidence supporting Shintani’s conjecture, the double sine function remains mysterious. To date, no non-trivial theoretical explanation has been found for the validity of Shintani’s conjecture.
The goal of this note is to generalize a key observation by Yamamoto [9], who showed that Shintani’s invariants $X(\mathfrak{f})$ can be expressed as a (limit of a) quotient of $q$-Pochhammer symbols. Specifically, we prove the following (Theorem 3.2):
Theorem 1.1.
Let $K=\mathbb{Q}(\sqrt{d})$ be a real quadratic number field with a totally positive unit $\varepsilon=\frac{a+b\sqrt{d}}{2}\in\mathcal{O}_{K,+}^{\times}$, where $a,b\in\mathbb{N}$, having a (minus) continued fraction expansion $\varepsilon=[[a]]$ of length one. Let $\mathfrak{f}=(u+v\sqrt{d})\in I_{K}$ be a principal ideal. Then there exist $g=g(\mathfrak{f})\in\mathbb{N}$ and $(x,y)\in\mathbb{Q}^{2}$, such that Shintani’s invariant $X(\mathfrak{f})$ is given by
(1.2) $X(\mathfrak{f})=\lim_{n\to\infty}\left|\frac{(x,y;\tau_{n-g})_{\infty}}{(x,y,\tau_{n+g})_{\infty}}\frac{(x,y;\tau_{-n-g})_{\infty}}{(x,y;\tau_{-n+g})_{\infty}}\right|,$
where
(1.3) $(x,y;\tau)_{\infty}=\prod_{k\geq 0}(1-e^{2\pi i(x\tau+y)e^{2\pi ik\tau}})$
is the $q$-Pochhammer symbol and $\tau_{n}=\frac{T_{n+1}(a)-ib\sqrt{d}}{T_{n}(a)}$.
The proof relies on a close link between the arithmetic of real quadratic fields and (minus) continued fractions [10, 8], which provides a natural description of the decomposition data $\{z_{k}\}_{k\in\{1,..,g\}}$ appearing in $X(\mathfrak{f})$. We also utilize an important product formula for the double sine function due to Shintani (Theorem 3.1), which enables telescoping expressions for Shintani’s invariants.
Several corollaries (3.1 -3.3) provide alternative formulations and refinements.
It is worth noting that the discrete parameter $\tau_{n}$, defined using classical Chebyshev polynomials $T_{n}(x)$, appears to be novel in this context.
In another direction, recent breakthroughs on Hilbert’s 12th problem from a $p$-adic perspective - see [4, 3] - highlight the importance of connecting Shintani’s archimedean viewpoint (via the double sine function or $q$-Pochhammer symbols) with these developments. Understanding this connection remains a significant challange. Let us mention a recent cohomological interpretation of Shintani invariants due to Kopp, see [5], and the possibly related works [1, 2].
We aim to present our results in as accessible a manner as possible, prioritizing clarity over generality to highlight the key ideas. In forthcoming work, we will further explore these invariants using the cyclic quantum dilogarithm.
Summary
In section 2), we recall Shintani’s definition of his invariants using the formalism of minus continued fractions (see [10, 8]), which is used to define the decomposition data. In section 3), after recalling important results due to Shintani and Yamamoto, we introduce a discretization of certain modular geodesics and use this to prove our main theorem expressing Shintani’s invariants in terms of $q$-Pochhammer symbols. In particular, we explain how one can express Shintani’s invariants using a single $q$ parameter. In the appendix we provide background on Chebyshev polynomials and prove a result concerning the size of the quantity $g(p)$, for $p$ a prime number.
Acknowledgements
We want to thank Giuseppe Ancona for useful feedback on the manuscript and Thomas Dreyfus for interesting discussions about $q$-difference equations. We also want to thank the anonymous referees for their valuable feedback.
1.1. Notations
Let $d$ be a positive, square-free integer and $K=\mathbb{Q}(\sqrt{d})$ be a real quadratic number field. We denote by $\mathcal{O}_{K}$ the ring of integers of $K$, by $\mathcal{O}_{K}^{\times}$ the group of units and by $\mathcal{O}_{K,+}^{\times}$ the group of totally positive units. The monoid of non-zero integral ideals of $\mathcal{O}_{K}$ is denoted by $I_{K}$. For any $x\in K$, we write $x^{\prime}$ for its Galois conjugate under the non-trivial automorphism $\sigma\in\mathrm{Gal(K/\mathbb{Q})}$.
We fix a totally positive unit $\varepsilon=\frac{a+b\sqrt{d}}{2}\in\mathcal{O}_{K,+}^{\times}$ with $a,b\in\mathbb{N}$. For an ideal $\mathfrak{f}\in I_{K}$, we define $g(\mathfrak{f})\in\mathbb{N}$ to be the smallest positive integer such that
(1.4) $\langle\varepsilon^{g(\mathfrak{f})}\rangle=\mathcal{O}_{K,+}^{\times}\cap(1+\mathfrak{f}),$
and we write $\varepsilon_{\mathfrak{f}}=\varepsilon^{g(\mathfrak{f})}$.
For $r\in\mathbb{R}$, we define $\langle r\rangle\in\mathbb{R}$ by $r-\langle r\rangle\in\mathbb{Z}$ and $0<\langle r\rangle\leq 1$. In the same way, we define $r]\in\mathbb{R}$ by $r-[r]\in\mathbb{Z}$ and $0\leq[r]<1$.
2. Shintani’s invariant
2.1. Shintani’s invariant
We follow the exposition in [7, 8]. Let $\mathfrak{C}\in\mathrm{Cl}_{K}(\mathfrak{f})$ be a (strict) ray class of conductor $\mathfrak{f}\in I_{K}$. Shintani defined the associated partial zeta function as
(2.1) $\zeta(s,\mathfrak{C})=\sum_{\mathfrak{a}\in\mathfrak{C},\mathfrak{a}\text{ integral}}\mathfrak{N}(\mathfrak{a})^{-s}$
and introduced the invariant
(2.2) $X(\mathfrak{C})=\mathrm{exp}(-\zeta^{\prime}(0,\mathfrak{C})+\zeta^{\prime}(0,\mathfrak{C}^{\prime})).$
These invariants $X(\mathfrak{C})$, known as the Shintani-Stark units, are conjectured to generate (essentially) the maximal abelian extension $K^{ab}$ of $K$. Shintani further expressed these invariants in terms of the double sine function:
(2.3) $\mathcal{S}(\omega,z)=\mathrm{exp}(\zeta_{2}^{\prime}(0,1+\omega-z,\omega)-\zeta^{\prime}_{2}(0,z,\omega)),$
where the Hurwitz zeta function $\zeta_{2}(s,z,(1,\omega))$ is given by
(2.4) $\zeta_{2}(s,z,\omega)=\sum_{n_{1},n_{2}=0}^{\infty}(n_{1}+n_{2}\omega+x)^{-s}.$
Shintani proved the following representation:
(2.5) $X(\mathfrak{C})=\prod_{k}\mathcal{S}(\varepsilon_{k},z_{k})\,\mathcal{S}(\varepsilon_{k}^{\prime},z_{k}^{\prime}),$
where $\mathcal{Z}_{\mathfrak{C}}=\{(\varepsilon_{k},z_{k})\}$ is a finite set known as the decomposition datum associated to the class $\mathfrak{C}$. Following [8], we define
(2.6) $X_{1}(\mathfrak{C})=\prod_{k}\mathcal{S}(\varepsilon_{k},z_{k})\text{ and }X_{2}(\mathfrak{C})=\prod_{k}\mathcal{S}(\varepsilon_{k}^{\prime},z_{k}^{\prime}).$
We will need the following identity (see Proposition 3.3.1 [8])
Proposition 2.1.
The double sine function satisfies
(2.7) $\mathcal{S}(\omega,z)=2\,\mathrm{sin}(\pi z)\,\mathcal{S}(\omega,z+\omega).$
To simplify the exposition, we restrict to the case where $\mathfrak{C}$ is the trivial class $1_{\mathfrak{f}}\in\mathrm{Cl}_{K}(\mathfrak{f})$. We write $X(\mathfrak{f})=X(1_{\mathfrak{f}})$ and $X_{i}(\mathfrak{f})=X_{i}(1_{\mathfrak{f}})$ for $i\in\{1,2\}$.
Furthermore, we assume that the generator $\varepsilon=\frac{a+b\sqrt{d}}{2}$ of $\mathcal{O}_{K,+}^{\times}$ has a minus continued fraction expansion (see [10]) of length one, i.e., we have
[…]
Under these assumptions, the partial zeta function admits the representation
(2.9) $\zeta(s,1_{\mathfrak{f}})=\sum_{\beta\in(\mu+\mathfrak{b})_{+}/\langle\varepsilon_{\mathfrak{f}}\rangle}\mathfrak{N}(\beta)^{-s},$
with the fractional ideal $\mathfrak{b}=\langle 1,\varepsilon\rangle$ and $\mu\in K^{\times}$ such that $\mathfrak{b}=(\mu)\,\mathfrak{f}$.
Remark 2.1.
If $\mathfrak{f}=(u+v\sqrt{d})\in I_{K}$ is a principal ideal, then we can take
(2.10) $\mu=\tfrac{1}{u+v\sqrt{d}}=\tfrac{u-v\sqrt{d}}{\mathfrak{N}(\mathfrak{f})}.$
In particular, if $v=0$, we have $\mu=\frac{1}{u}$.
By specialising Shintani’s cone decomposition theorem under our assumptions, we obtain the following result (cf., [8]):
Proposition 2.2.
Under the above assumptions, for each $k\in\mathbb{Z}$, there exists a unique pair $(x_{k},y_{k})\in\mathbb{Q}^{2}$ with $0<x_{k}\leq 1$, $0\leq y_{k}<1$, such that
(2.11) $x_{k}\varepsilon^{1-k}+y_{k}\varepsilon^{-k}\in\mu+\mathfrak{b}.$
In particular,
(2.12) $(\mu+\mathfrak{b})_{+}/\langle\varepsilon_{\mathfrak{f}}\rangle=\coprod_{k=1}^{g(\mathfrak{f})}\{(x_{k}+u)\varepsilon^{1-k}+(y_{k}+v)\varepsilon^{-k}|u,v\in\mathbb{Z}_{\geq 0}\}.$
This gives the decomposition datum
(2.13) $\mathcal{Z}_{\mathfrak{f}}=\{(\varepsilon,x_{k}\varepsilon+y_{k})\,|\,0\leq k\leq g(\mathfrak{f})-1\}$
of the class $1_{\mathfrak{f}}$, which leads to
(2.14) $X_{1}(\mathfrak{f})=\prod_{k=1}^{g(\mathfrak{f})}\mathcal{S}(\varepsilon,x_{k}\varepsilon+y_{k})\text{ and }X_{2}(\mathfrak{f})=\prod_{k=1}^{g(\mathfrak{f})}\mathcal{S}(\varepsilon^{\prime},x_{k}\varepsilon^{\prime}+y_{k}).$
Remark 2.2.
We sometimes abbreviate the decomposition datum as
(2.15) $\mathcal{Z}_{\mathfrak{f}}=\{(x_{k},y_{k})|k\in\{0,\dots,g(\mathfrak{f})-1\}\}$
for the decomposition datum.
Remark 2.3.
For an arbitrary real quadratic number field $K$, Lemma B.1 shows that for any prime $p\in\mathbb{N}$ not dividing the discriminant $\Delta$ of $K$, one always has
(2.16) $g(p)\,|\,p-\genfrac{(}{)}{}{}{\Delta}{p}$.
Remark 2.4.
Define $(\tilde{x}_{k},\tilde{y}_{k})=([x_{k}],y_{k})$ and set $\tilde{z}_{k}=\tilde{x}_{k}\varepsilon+y_{k}$. Then, the slightly modified invariant
(2.17) $\tilde{X}(\mathfrak{f})=\prod_{k=1}^{g(\mathfrak{f})}\mathcal{S}(\varepsilon,\tilde{z}_{k})\,\mathcal{S}(\varepsilon^{\prime},\tilde{z}_{k}^{\prime})$
satisfies (using Proposition 2.1)
(2.18) $\tilde{X}(\mathfrak{f})/X(\mathfrak{f})\in\mathbb{Q}^{\mathrm{ab}},$
hence
(2.19) $X(\mathfrak{f})\in K^{\mathrm{ab}}\Leftrightarrow\tilde{X}(\mathfrak{f})\in K^{\mathrm{ab}}.$
2.2. Explicit decomposition data for principal ideals
The matrix
(2.20) $U=\begin{bmatrix}a&-1\\ 1&0\end{bmatrix}\in\mathrm{SL}_{2}(\mathbb{Z})$
encodes the minus continued fraction expansion of $\varepsilon$, since
(2.21) $U\cdot\varepsilon=a-\frac{1}{\varepsilon}=\varepsilon,$
where the action is by Möbius transformation. This implies that the eigenvalues of $U$ are precisely $\{\varepsilon,\varepsilon^{\prime}\}$. Following [8], for $k\in\mathbb{Z}$ and an arbitrary $\mathfrak{f}\in I_{K}$, the decomposition data satisfy the recurrence:
(2.22) $(x_{k+1},y_{k+1})=(\langle ax_{k}+y_{k}\rangle,[-x_{k}])=(\langle ax_{k}-x_{k-1}\rangle,1-x_{k}),$
which corresponds to the matrix relation (modulo $1$)
(2.23) $\begin{bmatrix}x_{k+1}\\ y_{k+1}\end{bmatrix}=U^{T}\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}=\begin{bmatrix}a&1\\ -1&0\end{bmatrix}\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}$.
From Proposition 2.2, we thus obtain
Lemma 2.1.
For every $k\in\mathbb{Z}$, we have (upon properly taking modulo $1$)
(2.24) $(U^{T})^{g(\mathfrak{f})}\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}=\begin{bmatrix}x_{k}\\ y_{k}\end{bmatrix}$.
In particular,
(2.25) $(x_{0},y_{0})=(x_{g(\mathfrak{f})},y_{g(\mathfrak{f})}).$
To make this more explicit, we now assume that $\mathfrak{f}=(u+v\sqrt{d})\in I_{K}$ is a principal ideal. We set
(2.26) $\mu=\frac{u-v\sqrt{d}}{\mathfrak{N}(\mathfrak{f})}$
and write $\mu=x\varepsilon+y$ with $x,y\in\mathbb{Q}$. A straightforward calculation yields:
(2.27) $x=-\tfrac{2v}{b\,\mathfrak{N}(\mathfrak{f})}\text{ and }y=\tfrac{bu+av}{b\,\mathfrak{N}(\mathfrak{f})}.$
Thus, the initial pair $(x_{0},y_{0})$ for the cone decomposition is given by
(2.28) $(x_{0},y_{0})=(\langle x\rangle,[y]).$
For $k\geq 0$, the identity
(2.29) $(U^{T})^{k}=\begin{bmatrix}U_{k}(a)&U_{k-1}(a)\\ -U_{k-1}(a)&-U_{k-2}(a)\end{bmatrix}$,
(see formula (A.6)) leads to the explicit formula
(2.30) $(x_{k},y_{k})=(\langle U_{k}(a)x_{0}+U_{k-1}(a)y_{0}\rangle,[-U_{k-1}(a)x_{0}-U_{k-2}(a)y_{0}]).$
If we specialise further to $v=0$, we find
(2.31) $(x_{k},y_{k})=(\langle\tfrac{U_{k-1}(a)}{u}\rangle,[-\tfrac{U_{k-2}(a)}{u}]),$
so we only need to understand the behaviour of $U_{k}(a)$ modulo $u$. In this case, we have:
(2.32) $(x_{0},y_{0})=(1,\frac{1}{u}).$
2.3. Examples
We follow Shintani’s original examples from [7].
Example 2.1.
For $K=\mathbb{Q}(\sqrt{5})$, we have $\varepsilon=\tfrac{3+\sqrt{5}}{2}$. For the ideal $\mathfrak{f}=(4)$, we have $g=3$ and the decomposition data
(2.33) $\mathcal{Z}_{\mathfrak{f}}=\{(1,\tfrac{1}{4}),(\tfrac{1}{4},0),(\tfrac{3}{4},\tfrac{3}{4})\}.$
Thus,
(2.34) $X_{1}(4)=\mathcal{S}(\varepsilon,\varepsilon+\tfrac{1}{4})\mathcal{S}(\varepsilon,\tfrac{\varepsilon}{4})\mathcal{S}(\varepsilon,\tfrac{3\varepsilon}{4}+\tfrac{3}{4}),$
(2.35) $X_{2}(4)=\mathcal{S}(\varepsilon^{\prime},\varepsilon^{\prime}+\tfrac{1}{4})\mathcal{S}(\varepsilon^{\prime},\tfrac{\varepsilon^{\prime}}{4})\mathcal{S}(\varepsilon^{\prime},\tfrac{3\varepsilon^{\prime}}{4}+\tfrac{3}{4}).$
Shintani was able to explicitly compute
(2.36) $X(4)=\left(\tfrac{1+\sqrt{5}}{2}+\sqrt{\tfrac{1+\sqrt{5}}{2}}\right)^{-1}.$
Example 2.2.
Again, for $K=\mathbb{Q}(\sqrt{5})$, take $\mathfrak{f}=(4-\sqrt{5})$. Then $g=5$ and
(2.37) $\mathcal{Z}_{\mathfrak{f}}=\{(\tfrac{2}{11},\tfrac{1}{11}),(\tfrac{7}{11},\tfrac{9}{11}),(\tfrac{8}{11},\tfrac{4}{11}),(\tfrac{6}{11},\tfrac{3}{11}),(\tfrac{10}{11},\tfrac{5}{11})\}.$
Using different techniques (not based on the double sine function), Shintani showed:
(2.38) $X(4-\sqrt{5})=\frac{1}{2}\left(\tfrac{3+\sqrt{5}}{2}-\sqrt{\tfrac{3\sqrt{5}-1}{2}}\right).$
Example 2.3.
For $K=\mathbb{Q}(\sqrt{21})$, we have $\varepsilon=\tfrac{5+\sqrt{21}}{2}$. For $\mathfrak{f}=(3)$, we find $g=3$ and
(2.39) $\mathcal{Z}_{\mathfrak{f}}=\{(1,\tfrac{1}{3}),(\tfrac{1}{3},0),(\tfrac{2}{3},\tfrac{2}{3})\}.$
Shintani was able to av evaluate
(2.40) $X(3)=\frac{1}{2}\left(\tfrac{1+\sqrt{21}}{2}-\sqrt{\tfrac{3+\sqrt{21}}{2}}\right).$
Remark 2.5.
It is important to note that Shintani’s explicit evaluations in the above examples did not rely on properties of the double sine function $\mathcal{S}(\omega,z)$!
3. The $q$-Pochhammer function and Shintani’s invariant
3.1. Results of Shintani and Yamamoto
A lesser-known result due to Shintani [7] is the following product formula for the double sine function:
Theorem 3.1.
For $\mathrm{Im}(\tau)>0$, we have
(3.1) $\mathcal{S}(\tau,z)=i^{1/2}e^{\frac{\pi i}{12}(\tau+\frac{1}{\tau})}e^{\frac{\pi i}{2}(\frac{z^{2}}{\tau}-(1+\frac{1}{\tau})z)}\tfrac{\prod_{m\geq 0}(1-e^{2\pi i(m\tau+z)})}{\prod_{m\geq 1}(1-e^{2\pi i(\frac{-m+z}{\tau})})}.$
The special function appearing in Shintani’s theorem is well-known:
Definition 3.1.
The (infinite) $q$-Pochhammer symbol is defined by
(3.2) $(x,y;\tau)_{\infty}=\prod_{k\geq 0}(1-e^{2\pi i(k\tau+x\tau+y)})$
or, equivalently, in terms of $q=e^{2\pi i\tau}$ and $\zeta=e^{2\pi i(x\tau+y)}$, as
(3.3) $(\zeta;q)_{\infty}=\prod_{k\geq 0}(1-\zeta q^{k}).$
Lemma 3.1.
The $q$-Pochhammer symbol satisfies the following symmetries:
  • 1)
    $(x,y;\tau+1)_{\infty}=(x,x+y;\tau)_{\infty}$,
  • 2)
    $(x,y+1;\tau)_{\infty}=(x,y;\tau)_{\infty}$.
Later, Yamamoto [9] observed:
Proposition 3.1.
For $\varepsilon>0$, $x,y\in\mathbb{R}_{\geq 0}$ and $\mathrm{Im}(\tau)>0$, we have
(3.4) $\mathcal{S}(\varepsilon,x\varepsilon+y)=\lim_{\tau\to\varepsilon}|\mathcal{S}(\tau,x\tau+y)|=\lim_{\tau\to\varepsilon}\left|\tfrac{(x,y;\tau)_{\infty}}{(1-y,x;-\frac{1}{\tau})_{\infty}}\right|.$
Proof.
This follows from the definition (cf., (2.3)) and analyticity properties of the double sine function $\mathcal{S}(\omega,z)$ (see [7, 6].) . ∎
3.2. Modular geodesics
To make the limit $\tau\to\omega$ from Proposition 3.1 precise, we consider the modular geodesic connecting $\varepsilon$ and its Galois conjugate $\varepsilon^{\prime}=\varepsilon^{-1}$ in the upper half-plane $\mathbb{H}$. Define the matrix
(3.5) $A=\begin{bmatrix}1&1\\ \varepsilon^{\prime}&\varepsilon\end{bmatrix}$,
which diagonalizes $U$, i.e.,
(3.6) $A^{-1}UA=\begin{bmatrix}\varepsilon&0\\ 0&\varepsilon^{\prime}\end{bmatrix}$.
Then, the family
(3.7) $\tau_{t}=UA\begin{bmatrix}e^{t/2}&0\\ 0&e^{-t/2}\end{bmatrix}\cdot i$
describes the modular geodesic from $\varepsilon$ to $\varepsilon^{\prime}$, satisfying:
(3.8) $\mathrm{Im}(\tau_{t})>0\text{ for all }t\in\mathbb{R}\text{ and }\lim_{t\to\pm\infty}\tau_{t}=\varepsilon^{\pm 1}.$
Let us now discretize the variable $t$ by setting, for $n\in\mathbb{Z}$,
(3.9) $t_{n}=\mathrm{ln}(\varepsilon^{n}).$
This leads to
Lemma 3.2.
The sequence
(3.10) $\tau_{n}=\tau_{t_{n}}=\tfrac{T_{n+1}(a)+ib\sqrt{d}}{T_{n}(a)},$
for $n\in\mathbb{Z}$, gives a discrete approximation along the modular geodesic connecting $\varepsilon$ and $\varepsilon^{\prime}$, where $T_{n}(x)$ are the Chebyshev polynomials of the first kind.
Proof.
We have $UA=\begin{bmatrix}\varepsilon&\varepsilon^{\prime}\\ 1&1\end{bmatrix}$ and $\begin{bmatrix}\varepsilon^{n/2}&0\\ 0&\varepsilon^{-n/2}\end{bmatrix}\cdot i=i\varepsilon^{n}$, which leads to
(3.11) $\tau_{n}=\begin{bmatrix}\varepsilon&\varepsilon^{\prime}\\ 1&1\end{bmatrix}\cdot i\varepsilon^{n}=\tfrac{i\varepsilon^{n+1}+\varepsilon^{\prime}}{i\varepsilon^{n}+1}=\tfrac{(1-i\varepsilon^{n})(i\varepsilon^{n+1}+\varepsilon^{-1})}{\varepsilon^{n}T_{n}(a)}=\tfrac{T_{n+1}(a)+i(\varepsilon-\varepsilon^{\prime})}{T_{n}(a)}.$
Lemma 3.3.
For all $k\in\mathbb{Z}$, we have the shift relation
(3.12) $U^{k/2}\cdot\tau_{n}=\tau_{n+k}.$
Proof.
We have $\begin{bmatrix}\varepsilon^{1/2}&0\\ 0&\varepsilon^{-1/2}\end{bmatrix}=A^{-1}U^{1/2}A$, which implies $U^{k/2}A=A\begin{bmatrix}\varepsilon^{k/2}&0\\ 0&\varepsilon^{-k/2}\end{bmatrix}$. Thus
(3.13) $U^{k/2}\cdot\tau_{n}=U^{k/2}UA\begin{bmatrix}\varepsilon^{n/2}&0\\ 0&\varepsilon^{-n/2}\end{bmatrix}\cdot i=UA\begin{bmatrix}\varepsilon^{(n+k)/2}&0\\ 0&\varepsilon^{-(n+k)/2}\end{bmatrix}\cdot i=\tau_{n+k}.$
Remark 3.1.
Notice that $\lim_{n\to\pm\infty}\tfrac{T_{n+k}(a)}{T_{n}(a)}=\varepsilon^{\pm k}.$
3.3. A new formula for Shintani’s invariants
We now state our main result, which generalizes Yamamoto’s observation[9].
Theorem 3.2.
Under the previous assumptions, let $\mathfrak{f}=(u+v\sqrt{d})\in I_{K}$ be a principal ideal, with $g=g(\mathfrak{f})$, and let $(x_{0},y_{0})=(\langle x_{\mathfrak{f}}\rangle,[y_{\mathfrak{f}}])$ be the initial data (cf., (2.32)). Then:
(3.14) $X_{1}(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{(x_{0},y_{0};\tau_{n})_{\infty}}{(x_{0},y_{0};\tau_{n+2g})_{\infty}}\right|\text{ and }X_{2}(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{(x_{0},y_{0};\tau_{-n})_{\infty}}{(x_{0},y_{0};\tau_{-n+2g})_{\infty}}\right|.$
Proof.
The idea is to exploit Proposition 3.1, in order to obtain a description of the form
(3.15) $\mathcal{S}(\varepsilon,x_{k}\varepsilon+y_{k})=\lim_{n\to\infty}\left|\tfrac{(x_{k},y_{k};\tilde{\tau}_{k})_{\infty}}{(x_{k-1},y_{k-1};\tilde{\tau}_{k-1})_{\infty}}\right|.$
For this, we define
(3.16) $\tilde{\tau}_{k}=\tilde{\tau}_{n,k}=U^{g-k}\cdot\tau_{n}.$
Then, we have
(3.17) $\tilde{\tau}_{k-1}=U\cdot\tilde{\tau}_{k}=a-\frac{1}{\tilde{\tau}_{k}}\text{ \ \ \ and \ \ \ }\lim_{n\to\infty}\tilde{\tau}_{k}=\varepsilon,$
which means (using Proposition 3.1)
(3.18) $\mathcal{S}(\varepsilon,x_{k}\varepsilon+y_{k})=\lim_{n\to\infty}\left|\tfrac{(x_{k},y_{k};\tilde{\tau}_{k})_{\infty}}{(1-y_{k},x_{k};-\tfrac{1}{\tilde{\tau}_{k}})_{\infty}}\right|.$
Using Lemma 3.1 and the recursion $(x_{k},y_{k})=(\langle ax_{k-1}+y_{k-1}\rangle,1-x_{k-1})$ (see (2.22)), we obtain the following key observation for the denominator
(3.19) $(1-y_{k},x_{k};-\tfrac{1}{\tilde{\tau}_{k}})_{\infty}$ $=(1-y_{k},x_{k}-a(1-y_{k});a-\tfrac{1}{\tilde{\tau}_{k}})_{\infty}$
(3.20) $=(x_{k-1},\langle ax_{k-1}+y_{k-1}\rangle-ax_{k-1};\tilde{\tau}_{k-1})_{\infty}$
(3.21) $=(x_{k-1},y_{k-1};\tilde{\tau}_{k-1})_{\infty}.$
Using $(x_{0},y_{0})=(x_{g},y_{g})$, $U^{g}\cdot\tilde{\tau}_{g}=\tilde{\tau}_{0}$ and $\tilde{\tau}_{g}=\tau_{n}$, we get
(3.22) $X_{1}(\mathfrak{f})$ $=\prod_{k=1}^{g}\mathcal{S}(\varepsilon,x_{k}\varepsilon+y_{k})=\lim_{n\to\infty}\left|\prod_{k=1}^{g}\tfrac{(x_{k},y_{k};\tilde{\tau}_{k})_{\infty}}{(x_{k-1},y_{k-1};\tilde{\tau}_{k-1})_{\infty}}\right|$
(3.23) $=\lim_{n\to\infty}\left|\tfrac{(x_{g},y_{g};\tilde{\tau}_{g})_{\infty}}{f(x_{0},y_{0};\tilde{\tau}_{0})_{\infty}}\right|=\lim_{n\to\infty}\left|\tfrac{(x_{g},y_{g};\tilde{\tau}_{g})_{\infty}}{(x_{g},y_{g};U^{g}\cdot\tilde{\tau}_{g})_{\infty}}\right|$
(3.24) $=\lim_{n\to\infty}\left|\tfrac{(x_{0},y_{0};\tau_{n})_{\infty}}{(x_{0},y_{0};\tau_{n+2g})_{\infty}}\right|.$
Moreover, setting
(3.25) $\tilde{\tau}_{k}^{\prime}=\tilde{\tau}_{n,k}^{\prime}=U^{g-k}\cdot\tau_{-n},$
we have
(3.26) $\tilde{\tau}_{k-1}^{\prime}=U\cdot\tilde{\tau}_{k}^{\prime}\ \ \ \text{ and }\ \ \ \lim_{n\to\infty}\tilde{\tau}_{k}^{\prime}=\varepsilon^{\prime},$
showing with same argument (verbatim)
(3.27) $X_{2}(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{(x_{0},y_{0};\tau_{-n})_{\infty}}{(x_{0},y_{0};\tau_{-n+2g})_{\infty}}\right|.$
Remark 3.2.
A general version of this theorem, for general $\mathfrak{f}$ and $\mathfrak{C}$, is available. For clarity, we only consider the case where $\mathfrak{f}$ is a principal ideal.
Corollary 3.1.
With previous assumptions, for all $k\in\{0,\dots,g-1\}$, the invariant satisfies
(3.28) $X(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{(x_{k},y_{k};\tau_{n})_{\infty}(x_{k},y_{k};\tau_{-n})_{\infty}}{(x_{k},y_{k};\tau_{n+2g})_{\infty}(x_{k},y_{k};\tau_{-n+2g})_{\infty}}\right|=\lim_{n\to\infty}\left|\tfrac{(x_{k},y_{k};\tau_{n-g})_{\infty}(x_{k},y_{k};\tau_{-n-g})_{\infty}}{(x_{k},y_{k};\tau_{n+g})_{\infty}(x_{k},y_{k};\tau_{-n+g})_{\infty}}\right|.$
Proof.
The first equality simply comes from the shift invariance $(x_{k},y_{k})\mapsto(x_{k+1},y_{k+1})$ of the Shintani invariant $X(\mathfrak{f})$. The second equality comes from the shift $n\mapsto n-g$. ∎
Corollary 3.2.
With the previous assumptions, if $\mathfrak{f}=(u)\in I_{K}$, and we define $q_{n}=e^{2\pi i\tau_{n}}$, $\zeta_{u}=e^{2\pi i/u}$, then:
(3.29) $X(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{(\zeta_{u};q_{n})_{\infty}(\zeta_{u};q_{-n})_{\infty}}{(\zeta_{u};q_{n+2g})_{\infty}(\zeta_{u};q_{-n+2g})_{\infty}}\right|=\lim_{n\to\infty}\left|\tfrac{(\zeta_{u};q_{n-g})_{\infty}(\zeta_{u};q_{-n-g})_{\infty}}{(\zeta_{u};q_{n+g})_{\infty}(\zeta_{u};q_{-n+g})_{\infty}}\right|.$
Proof.
This follows from $(x_{0},y_{0})=(1,\tfrac{1}{u})$, see (2.32). ∎
The formulas (3.14), (3.28) and (3.29) contain two different $q$ parameters. We want to show that we can express $X(\mathfrak{f})$ in terms of a single $q$ parameter:
Corollary 3.3.
Continuing with the same assumptions, let $\mathfrak{q}=e^{-2\pi\sqrt{d}}$, $\zeta_{0}=e^{2\pi i(x_{0}\varepsilon+y_{0})}$, $T_{n}=T_{n}(a)$ and
(3.30) $\zeta_{n,r}=\zeta_{0}e^{2\pi ir\frac{T_{n+1}}{T_{n}}}e^{-2\pi b\sqrt{d}\frac{r}{T_{n}}},$
for $0\leq r<T_{n}$. Then, we have
(3.31) $X_{1}(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{\prod_{r=0}^{T_{n-g}}(\zeta_{n-g,r};\mathfrak{q})_{\infty}}{\prod_{r=0}^{T_{n+g}}(\zeta_{n+g,r};\mathfrak{q})_{\infty}}\right|\text{ and }X_{2}(\mathfrak{f})=\lim_{n\to\infty}\left|\tfrac{\prod_{r=0}^{T_{n+g}}(\zeta_{-n-g,r};\mathfrak{q})_{\infty}}{\prod_{r=0}^{T_{n-g}}(\zeta_{-n+g,r};\mathfrak{q})_{\infty}}\right|.$
Proof.
This follows from the multiplicative dependence of $q_{n-g}=e^{2\pi i\tau_{n-g}}$ and $q_{n+g}=e^{2\pi i\tau_{n+g}}$, as
(3.32) $\mathfrak{q}=q_{n-g}^{T_{n-g}}=q_{n+g}^{T_{n+g}}.$
From
(3.33) $(\zeta_{0};q_{n})_{\infty}=\prod_{k\geq 0}(1-\zeta_{0}q_{n}^{k})=\prod_{s\geq 0}\prod_{r=0}^{T_{n}-1}(1-\zeta_{0}q_{n}^{sT_{n}}q_{n}^{r})=\prod_{s\geq 0}\prod_{r=0}^{T_{n}-1}(1-\mathfrak{q}^{s}\zeta_{n,r})=\prod_{r=0}^{T_{n}-1}(\zeta_{n,r};\mathfrak{q})_{\infty},$
where $k=sT_{n}+r$, with $0\leq r<T_{n}$, the result follows. ∎
3.4. Examples
The examples from section 2.3 can now be rewritten:
Example 3.1.
For $K=\mathbb{Q}(\sqrt{5})$ and $\mathfrak{f}=(4)$, we have $a=3$, $g=3$ and
(3.34) $X(4)=\lim_{n\to\infty}\left|\tfrac{(\zeta_{4};q_{n-3})_{\infty}(\zeta_{4};q_{-n-3})_{\infty}}{\zeta_{4};q_{n+3})_{\infty}(\zeta_{4};q_{-n+3})_{\infty}}\right|.$
Example 3.2.
Again, for $K=\mathbb{Q}(\sqrt{5})$, $\mathfrak{f}=(4-\sqrt{5})$, $g=5$ and using $(x_{0},y_{0})=(2/11,1/11)$, we get:
(3.35) $X(4-\sqrt{5})=\lim_{n\to\infty}\left|\tfrac{(2/11,1/11;q_{n-5})_{\infty}(2/11,1/11;q_{-n-5})_{\infty}}{2/11,1/11;q_{n+5})_{\infty}(2/11,1/11;q_{-n+5})_{\infty}}\right|.$
Example 3.3.
For $K=\mathbb{Q}(\sqrt{21})$, $\mathfrak{f}=(3)$, $a=5$ and $g=3$, we obtain:
(3.36) $X(3)=\lim_{n\to\infty}\left|\tfrac{(\zeta_{3};q_{n-3})_{\infty}(\zeta_{3};q_{-n-3})_{\infty}}{\zeta_{3};q_{n+3})_{\infty}(\zeta_{3};q_{-n+3})_{\infty}}\right|.$
Remark 3.3.
It remains a highly nontrivial problem to recover Shintani’s explicit evaluations of the above examples purely from the properties of the $q$-Pochhammer symbol.
Appendix A Chebyshev polynomials of the first and second kind
For $n\in\mathbb{Z}$, the (normalized) $n$-th Chebyshev polynomial of the first kind $T_{n}(x)\in\mathbb{Z}[x]$ is defined by the identity
(A.1) $T_{n}(x+x^{-1})=x^{n}+x^{-n}.$
Similarly, the (normalized) $n$-th Chebyshev polynomial of the second kind $U_{n}(x)\in\mathbb{Z}[x]$ is defined via:
(A.2) $U_{n-1}(x+x^{-1})=\tfrac{x^{n}-x^{-n}}{x-x^{-1}}.$
In particular, we see
(A.3) $T_{n}(a)=T_{n}(\varepsilon+\varepsilon^{-1})=\varepsilon^{n}+\varepsilon^{-n}\in\mathbb{N},$
(A.4) $U_{n}(a)=U_{n}(\varepsilon+\varepsilon^{-1})=\tfrac{\varepsilon^{n}-\varepsilon^{-n}}{\varepsilon-\varepsilon^{-1}}\in\mathbb{N}.$
These polynomials satisfy the standard identities:
  • Multiplicative identities:
    (A.5) $T_{n}(x)T_{m}(x)=T_{n+m}(x)+T_{n-m}(x),\forall n,m\in\mathbb{Z},$
    (A.6) $U_{n}(x)U_{m}(x)=U_{n+m}(x)+U_{n-m}(x),\forall n,m\in\mathbb{Z}.$
  • Recurrence relations:
    (A.7) $T_{n+1}(x)=xT_{n}(x)-T_{n-1}(x),\forall n\in\mathbb{Z},$
    (A.8) $U_{n+1}(x)=xU_{n}(x)-U_{n-1}(x),\forall n\in\mathbb{Z}.$
  • Symmetry:
    (A.9) $T_{n}(x)=T_{|n|}(x),\forall n\in\mathbb{Z},$
    (A.10) $U_{-1+n}(x)=-U_{-1-n}(x),\forall n\in\mathbb{Z}_{\geq 0}.$
Example A.1.
We have
  • $T_{-2}(x)=x^{2}-2$, $T_{-1}(x)=x$, $T_{0}(x)=2$, $T_{1}(x)=x$, $T_{2}(x)=x^{2}-2$, $T_{3}(x)=x^{3}-3x$,
  • $U_{-2}(x)=-1$, $U_{-1}(x)=0$, $U_{0}(x)=1$, $U_{1}(x)=x$, $U_{2}(x)=x^{2}-1$, $U_{3}(x)=x^{3}-2x$.
Appendix B A result about $g(p)$
Lemma B.1.
Let $p\in\mathbb{N}$ be a prime number not dividing the discriminant $\Delta$ of $K$. Then, with the usual notation for the Jacobi symbol, we have:
(B.1) $g(p)\ |\ p-\genfrac{(}{)}{}{}{\Delta}{p}$.
Proof.
Recall $\varepsilon=\tfrac{a+b\sqrt{d}}{2}$, with $a,b\in\mathbb{N}$, and set $l=p-\genfrac{(}{)}{}{}{\Delta}{p}$. When we write $e\equiv f$, we always mean $e\equiv f\text{ mod $p$}$.
Let us start with the case $(p)$ is non-split, i.e., in particular $\genfrac{(}{)}{}{}{\Delta}{p}=-1$. We need to show that
(B.2) $p\ |\ 2^{p+1}(\varepsilon^{p+1}-1).$
As
(B.3) $\binom{p}{k}\equiv\left\{\begin{array}[]{c}0\text{ , if $0<k<p$}\\ \ 1\text{ , if $k\in\{0,p\}$}\end{array}\right$.,
we have
(B.4) $2^{p+1}\varepsilon^{p+1}=(a+b\sqrt{d})^{p}(a+b\sqrt{d})=a^{p+1}+b^{p+1}d^{(p+1)/2}+(a^{p}b+ab^{p}d^{(p-1)/2})\sqrt{d}.$
From the assumptions, we know
(B.5) $d^{(p+1)/2}\equiv-d\text{ and }d^{(p-1)/2}\equiv-1.$
Moreover, using Fermat’s little theorem, we see
(B.6) $a^{p+1}+b^{p+1}d^{(p+1)/2}\equiv a^{2}-db^{2}\equiv 4$
and
(B.7) $a^{p}b+ab^{p}d^{(p-1)/2}\equiv ab-ab=0.$
This gives
(B.8) $2^{p+1}\varepsilon^{p+1}\equiv 4$
and thus $p\ |\ 2^{p+1}(\varepsilon^{p+1}-1)$.
Now we look at the case where $(p)$ splits. In this case we need to show that
(B.9) $p\ |\ (\varepsilon^{l}-1)(\varepsilon^{-l}-1)=2-(\varepsilon^{l}+\varepsilon^{-l})=2-T_{l}(a).$
We calculate
(B.10) $2^{l}\,T_{l}(a)=2\sum_{k=0}^{l/2}\binom{l}{2k}a^{2k}(b^{2}d)^{l/2-k}.$
If $\genfrac{(}{)}{}{}{\Delta}{p}=1$, there exists $1\leq r\leq l/2$ such that $r^{2}\equiv d$. Using $\binom{p-1}{2k}\equiv 1$ for $0\leq k\leq l/2$ and Fermat’s little theorem, we obtain
(B.11) $2^{l}\,T_{l}(a)\equiv T_{l}(a)\equiv 2(br)^{l}\sum_{k=0}^{l/2}(a^{2}b^{-2}d^{-1})^{k}\equiv 2\tfrac{(a^{2}b^{-2}d^{-1})^{l+1}-1}{(a^{2}b^{-2}d^{-1})-1}\equiv 2\tfrac{(a^{2}b^{-2}d^{-1})-1}{(a^{2}b^{-2}d^{-1})-1}=2.$
If $\genfrac{(}{)}{}{}{\Delta}{p}=-1$, we have again $d^{l/2}\equiv-d$. From
(B.12) $\binom{p+1}{k}\equiv\left\{\begin{array}[]{c}0\text{ ,if $2\leq k\leq p-1$}\\ \ 1\text{, if $k\in\{0,1,p,p+1\}$}\end{array}\right$.,
together with Fermat’s little theorem, we get
(B.13) $2^{l}\,T_{l}(a)\equiv 2(a^{l}+b^{l}d^{l/2})\equiv 2(a^{2}-db^{2})=8.$
Thus, we get $p\ |\ T_{l}(a)-2$, finishing the proof. ∎